Hunt Taxidermy, Inc. is concerned about the taxes paid by the company in 2018. In addition to $12.7 million of taxable income, the firm received $1,260,000 of interest on state-issued bonds and $1,000,000 of dividends on common stock it owns in Oakdale Fashions, Inc. (Use Table 2.3.) Calculate Hunt Taxidermys tax liability. (Enter your answer in dollars not in millions. Round your answer to the nearest dollar amount.) Answer is complete and correct. Tax liability |$ 4.450.000
Calculate Hunt Taxidermys average tax rate. (Round your answer to 2 decimal places.) Answer is complete but not entirely correct. Average tax rate 29 29.75 x%
table 2.3 Corporate Tax Rates as of 2015 Pay this AmountPlus this Percentage on on Base Income Taxable In come Anything Over the Base O-$50,000 50,001-$75,000 75,001-5100,000 100,001-5335,000 335,001-$10,000,000 10,000,001-$15,000,000 15,000,001-518,333 333 Over $18,333,333 0 7,500 13,750 22,250 13,900 3,400,000 5,150,000 6,416,667 15% 25 34 39 34 35 38 35
Answer #1

Average tax rate = Tax payable/Taxable income =

= 4450000/13000000 = 34.23%

Taxable income = 12700000+30% of 1000000 = 13,000,000